A ray of light is incidented at angle 60° in air on air glass plane surface.find the angle of refraction of the glass (aΠg =1•5) in calculation method


1
Expert's answer
2021-07-12T12:13:12-0400

The angle of refraction is given by the Snell's law:



"\sin i\cdot n_1 = \sin r\cdot n_2"

where "i = 60\degree" is the angle of incidence, "r" is the angle of refraction, and "n_1 = 1, n_2 = 1.5" are the indices of refraction of air and glass respectively.

Thus, obtain:


"r = \arcsin\left( \sin i\cdot \dfrac{n_1}{n_2} \right) = \arcsin\left( \sin60\degree\cdot \dfrac{1}{1.5} \right) \approx 35\degree"

Answer. "35\degree".